OpenStudy (anonymous):

what is a counter example of x>1/x

OpenStudy (anonymous):

is it\[x<\frac{ 1 }{ x } ?\]

OpenStudy (anonymous):

thank you , I have tru=ouble figuring out counter examples to prove a statement is false

OpenStudy (anonymous):

I need to prove that this equation is not always correct

OpenStudy (anonymous):

mathematical Excursions

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf x< \frac{1}{x} }\) \(\Huge\color{blue}{ \sf 2< \frac{1}{2} }\) well, 1/2 is less than 2, not greater than 2. \(\Huge\color{green}{ \sf x> \frac{1}{x} }\) \(\Huge\color{green}{ \sf .5> \frac{1}{.5} }\) ... \(\Huge\color{green}{ \sf 1/2> 2 }\)

OpenStudy (anonymous):

When you explain it like that it makes sense. Thank you Solomon

OpenStudy (solomonzelman):

You welcome :)

OpenStudy (anonymous):

i did not follow your question. let x>0 \[x>\frac{ 1 }{ x },x^2>1\] \[\left| x \right|>1,x<-1(rejected) ~and~x>1\] it is false in 0<x<1

OpenStudy (anonymous):

let x<0 \[x>\frac{ 1 }{ x },x^2<1,\left| x \right|<1,-1<x<1\] it is true in -1<x<0 for all other negative values it is false.

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

yw